an invested company pays 5% interest compounded semiannually. you want to have $16000 in the future, how much should you deposit now to have that amount 5 years from now?
Thank you

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

an invested company pays 5% interest compounded semiannually. you want to have $16000 in the future, how much should you deposit now to have that amount 5 years from now?
--------------------------------
A(t) = P(1+(r/n))^(nt)
----
16000 = P(1+(0.05/2))^(2*5)
---
16000 = P(1.28)
----
P = $12,499.17
========================
cheers,
Stan H.

Answer by checkley77(12844)

x(1+.05/2)^5*2=16,000
x(1+.025)^10=16,000
x(1.025)^10=16,000
x*1.28=16,000
x=16,000/1.28
x=$12,500 is the necessary investment to have $16,000 in 5 years.